Consider a series circuit including a 4Ω, a 16mH-inductor, and a power supply of 9 volts.

How many times constant does it take for the current to reach 99.9% of its final value?


1
Expert's answer
2021-12-19T18:23:11-0500

R=4 "\Omega"

L=16mH

V=9V

"I={0.999I_0}"

"\tau=\frac{L}{R}"

"\tau=\frac{16\times10^3}{4}=4\times10^{-3}" Sec

"I=I_0(1-e^{(\frac{t}{\tau})})"

"0.999I_0 =I_0(1-e^{(\frac{-t}{\tau})})"

"0.999=(1-e^{(\frac{-t}{\tau})})"

"1\times10^{-3}=e^{(\frac{-t}{\tau})}"

"10^3=e^{\frac{t}{\tau}}"

"ln1000=\frac{t}{\tau}"

"t=\tau ln1000"

"t=27.63\times10^{-3}=27.63msec" "t=27.63 msec" times constant does it take for the current to reach 99.9% of its final value